I have turned to you for help in solving this and received an answer, but my teacher has advised me that it is not in correct format. Can anyone please help me make this correct format. Thank you so much:-)
2(x-5)=3
Answer is: X1=3.77 and X2=6.22,
How do I make this look like an algebra equation for an answer?

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

2(x-5)=3
2x-10=3
2x=13
x=13/2
Cheers,
Stan H.

Answer by checkley71(8403)

DON'T KNOW WHAT YOU DID BUT YOUR ANSWERS ARE NOT CORRECT.
TRY THE FOLLOWING METHOD:
2(X-5)=3
2X-10=3
2X=3+10
2X=13
X=13/2 OR X=6.5
PROOF
2(6.5-5)=3
2(1.5)=3
3=3